OpenStudy (anonymous):

Factor this expression completely, then place the factors in the proper location. 1/3x^3-1/27y^3

OpenStudy (anonymous):

\[a = \sqrt[3]{\frac{1}{3}}\] and \[b = \frac{1}{3}\]

OpenStudy (anonymous):

\[(xa - yb)(x^2a^2 + y^2b^2 + xyab) \]

OpenStudy (anonymous):

a^3 - b^2 = (a -b)(a^2+b^2 + ab)

OpenStudy (anonymous):

how did you create the division line?

OpenStudy (anonymous):

{x over y}